# Detection of a Type-C QPO during the soft-to-hard transition in Swift J1727.8 –1613

### v1

@@ 3.2 The Real and Imaginary parts of the cross spectrum
Following Méndez et al. (2024), the Real and Imaginary parts of the CS were fitted in a similar way as the PS, where they are decomposed into a linear combination of Lorentzians multiplied by the phase lag dependency using the relation:
@@
with $C=\sqrt{AB}$ and the frequency-dependent phase lags between the two signals is defined as:
@@
Here, we explicitly treat the phase lags of the i-th Lorentzian as a function of the parameters $p_{\textit{j,i}}$. In the rest of the paper, we assume the phase lags of each Lorentzian to be constant with frequency: $g_{i}(\nu;p_{j,i})=\Delta\phi_{\textit{i}}$, where $\Delta\phi_{\textit{i}}$ is a parameter of the model. This is the constant phase lag model as defined in Méndez et al. (2024). The phase lags of J1727 are close to zero across the Fourier frequencies, thus, the amplitude of the Imaginary part of the CS is smaller than the Real part. Consequently, the cross vectors can be rotated by 45° to have comparable Real and Imaginary parts, enhancing more stable fits in Xspec, without impacting the best-fitting parameters (c.f. Méndez et al., 2024). Therefore, by multiplying the cross vector (decomposed into a Real and Imaginary part) by the rotation matrix, we calculated the Real and Imaginary parts of the CS as $\mathrm{\Re[\cos{(\pi/4)}]\ -\ \Im[\sin{(\pi/4)}]}$ and $\mathrm{\Re[\sin{(\pi/4)}]\ +\ \Im[\cos{(\pi/4)}]}$, respectively. The central frequencies and FWHMs of the three Lorentzians in the CS are tied to those in the 0.3–2 keV band of the PS. Meanwhile, the normalization of the Real part of the CS is given by the parameter $C$, previously defined. Table 1 shows the best-fitting parameters of the phase lags of the CS between 0.3–2 keV and 2–12 keV. The fitted rotated Real (green) and rotated Imaginary (pink) parts of the CS are shown on the right panel of Fig. 1. Below 1.5 Hz, the rotated Real and Imaginary parts of the CS are overlapping in the two energy bands, for both observations. Above 1.5 Hz, the rotated Imaginary part of the CS in the 2–12 keV energy range is higher than the rotated Real part. A bump is observed in the rotated Imaginary part of the CS at $\sim$3 Hz, and is well-fitted by a narrow Lorentzian. In both observations, we identify a narrow Lorentzian in the PS and the CS with a stronger Imaginary component compared to the Real part. We name this QPO the “Imaginary QPO”, as first expressed by Méndez et al. (2024). While the rms of the Imaginary QPO stays more or less constant when the system hardens, the Real part’s increases.
Figure 2: The phase lags in radian (top) and the coherence function (bottom) for the second (left panel) and third (right panel) EPIC-PN observation of Swift J1727.8–1613. The continuous lines show the predicted model from the fits of the PS and CS represented in Fig. 1. The bottom panels show the residuals defined as (data-model)/error. The red dotted vertical line indicates the central frequency of the Imaginary QPO for each observation.
@@ 3.3 The phase lags and coherence
The last column of Table 1 shows the best-fitting parameters for the phase lags. The phase lags of the three Lorentzians increase by $\sim$37%, $\sim$13%, and $\sim$19(for L1, L2, and L3, respectively), between the second and third observation. From the PS and CS fits, we derive a model of the phase lags and the coherence, calculated in the same energy bands as the power and cross spectra. Fig. 2 displays the phase lags (top panel) and the coherence (bottom panel) for the two observations, together with the predicted model. Below 1 Hz, the phase lags are on average constant around $\sim$0 rad, with a small excess around $\sim$0.2 Hz and $\sim$0.5 Hz, significant at the 68% confidence level. Then, the phase lags reach a maximum value at the Imaginary QPO frequency, marked by the vertical dotted red line, and stabilize to a higher constant value around $\sim$1.8 rad after that, for both observations. However, in the second observation, the phase lags are narrower and are dominated by the uncertainties above 7 Hz for both observations.
While the phase lags increase at the Imaginary QPO’s central frequency, the coherence drops. Below 1 Hz, the average coherence value is around $\sim$1, it then drops to $\sim$0.8 at the frequency of the “Imaginary QPO” and remains constant around $\sim$0.9 after the dip. Similarly to the phase lags, the residuals drop significantly around $\sim$0.5 Hz, aligning with a slight decrease in the coherence. These features suggest the presence of an additional component contributing to the variability at low frequency. However, we did not add another Lorentzian to account for this component in our model, as we aim to fit the prominent variable components with as few Lorentzians as possible.
@@ 3.4 The energy dependence of the QPO lags and rms amplitude
To analyze the QPO lags and rms amplitude energy dependence, we computed the Fast Fourier Transforms (FFTs) and CS for each narrow energy band relative to the full (0.3–12 keV) energy band, at the Imaginary QPO frequency. The data were then fitted using the multi-Lorentzian model previously defined, with central frequencies and FWHMs fixed to the best fit values listed in Table 1. Fig. 3 shows the energy dependence of the fractional rms (top) and the phase lags (bottom).
@@
On the other hand, the phase lags reach a minimum value at 1 keV. In Epoch XMM2, the phase lags decrease from $\sim$ –0.08 rad to $\sim$ –0.28 rad between 0.5 keV and 1 keV and increase again from $\sim$ –0.21 rad at 1.2 keV to $\sim$0.58 rad at 8 keV. In Epoch XMM3, a similar behaviour is observed with a decrease from $\sim$ –0.09 rad at 0.5 keV to a minimum value of $\sim$ –0.31 rad at 1 keV, increasing above 1 keV.
@@ 4 Discussion and conclusions
In section 3, we showed that the XMM EPIC-PN data of the BH LMXB Swift J1727.8–1613 reveal the presence of a Type-C QPO in the soft–to–hard transition. While the characteristic Lorentzians’ frequencies shift to lower values as the source hardens, the phase lags increase and the coherence function drops at the QPO frequency, reinforcing the possible cause–to–effect relationship between the Type-C QPO and the features observed in the lags and the coherence. An in-depth discussion is provided below.
@@ 4.1 The Type-C QPO
We identify a narrow Lorentzian in the PS characterized by a higher rms amplitude in the Imaginary part than in the Real part of the CS, where the Real counterpart’s rms increases between the second and the third observation. The origin of the Imaginary QPO is still unclear: In Cyg X-1, König et al. (2024) suggested a beat between the two broad Lorentzians, which could create a third component causing the QPO, often observed in the PS, whereas Méndez et al. (2024) argue that it comes from an independent phenomenon. Nonetheless, the Imaginary QPO of the source seems to correspond to the QPO observed in the PS.
@@
The QPO identified in J1727’s PS and CS has a central frequency at $\sim$2.6 Hz in the second observation, shifting to $\sim$2 Hz in the third observation. Furthermore, the total rms is $\sim$17and $\sim$18in the 0.3–2 keV energy band, for the second and last observation, respectively, reaching $\sim$30in the 2–12 keV energy band for both observations. Given J1727’s brightness and high-rms LFQPO in both the PS and CS, we classify the LFQPO, observed in the soft–to–hard transition, as a Type-C QPO (Motta et al., 2011). In the rms–intensity diagram for black hole transients defined by Muñoz-Darias et al. (2011), the source would be close to the hard line, the designated position of Type-C QPOs. However, the $Q$ factor values for J1727 are lower than those usually observed for Type-C QPOs ($\sim$7–10). Low $Q$ factors in Type-C QPOs were also observed in MAXI J1348–630 (Alabarta et al., 2022) and MAXI J1820+070 (Ma et al., 2023), ranging between 0.2 and 7.
@@
On the other hand, as the system hardens, the QPO moves to lower frequencies. This frequency shift – probably induced by the Comptonizing plasma dominating the emission in the hard state – has previously been observed in other XBs (MAXI J1820+070; Méndez et al., 2024; Bellavita et al., 2025, and Cyg X-1; König et al., 2024; Fogantini et al., 2025). Furthermore, during the soft–to–hard transition of MAXI J1348–630, Alabarta et al. (2025) interpreted the QPO frequency shifting to lower frequencies as indicative of coronal expansion. In the case of GRS 1915+105, Méndez et al. (2022) and García et al. (2022) explain the frequency shift by a perpendicular contraction of the corona with respect to the accretion disk. Thus, the frequency shift suggests a geometric change in the corona.
@@
Additionally, observed in the soft–to–hard transition, the Type-C QPO can be attributed to the expansion of the truncated disk inner radius or by feedback processes from the corona (Bellavita et al., 2022). The Lense-Thirring precession, due to the inner accretion flow precessing with respect to the thermal disk (see Ingram et al., 2009), has also been proposed as the origin of Type-C QPOs in XBs, supported by studies showing a tight anti-correlation between the outer radius of the Comptonizing region and $\mathrm{\nu_{QPO}}$ (Kubota et al., 2024). However, the analysis of GRS 1915+105 by Nathan et al. (2022) reveals an inner disk radius too small and a reprocessing time too large to support this model.
@@
Apart from the Type-C QPO, the broadband Lorentzians, L1 and L3, similarly shift to lower frequencies as the source hardens. However, in the 0.3–2 keV energy band, as the source transitions, their rms amplitude increases, unlike that of the Type-C QPO. This behaviour was previously observed in GX 339-4 (Muñoz-Darias et al., 2011). In the 2–12 keV hard band, for a harder emission, the rms increases for L3 but decreases for L1. Thus, as suggested by Bellavita et al. (2025), the two broad Lorentzians seem independent from each other and might come from independent variability processes than the QPO’s. König et al. (2024) support this assumption proposing L1 to come from the accretion disk and modulated by the Comptonization, and L3 to be associated with the Comptonized medium. On the other hand, Wilkinson & Uttley (2009) link the L1 variability to an unstable accretion flow in the disk (though they reached that conclusion by fitting the covariance spectrum with the same model as the time-averaged one, which only holds if only the overall normalization of the model varies).
4.2 Phase lags jump
@@
At the QPO frequency, the phase lags reach a maximum, also called the “cliff”, and was previously observed in other XBs (Méndez et al., 2024; Bellavita et al., 2025). Yet, before and after the QPO frequency, the phase lags remain constant around $\sim$0 rad and $\sim$0.18 rad, respectively, for both observations, with a sharper “cliff” in the third observation. In the frequency range in which one Lorentzian dominates the PS, the phase lags remain on average constant, where each Lorentzian has its own lags, potentially due to separate resonances (Méndez et al., 2024). On the contrary, at the intersection of the two Lorentzians, the phase lags transition from one constant value to another one. Nowak et al. (1999b) attributed this difference to the total phase lags at each Fourier frequency being equal to the average phase lags of the individual Lorentzians, weighted by the product of their amplitudes in the two energy bands at which they are calculated (defined as the small-angle approximation of equation 9 in Méndez et al., 2024). As a result, the phase lags are constant over a given frequency range when the Fourier amplitudes of one Lorentzian dominate.
@@
Besides, the “cliff” is not observed in the hard energy bands. Similarly observed in the rising phase of the outburst of J1727, Yu et al. (2024) did not find the phase lags jump in the 1–10 keV and 25–150 keV energy bands, but still detected it in the soft energy bands, 1–2 keV and 2–10 keV, this may suggest that soft energies are essential to observe the phase-lag jump, regardless of the transition direction. Perhaps indicating that the QPO may originate from variations in a soft component linked to the accretion disk. Furthermore, as the source hardens, the Comptonization-driven variability from the corona intensifies and may introduce additional variability in the accretion disk.
@@ 4.3 Coherence drop
Overall, the phase-lags jump and the coherence drop seem to be generated by low-energy variations from a soft component in the accretion disk at the Type-C QPO frequency, induced by variations in the corona geometry (Méndez et al., 2022; Alabarta et al., 2025). Only observed in the decaying phase of the outburst (Méndez et al., 2024; Alabarta et al., 2025; König et al., 2024; Bellavita et al., 2025; Fogantini et al., 2025), these features are significant at low X-ray flux in the soft–to–hard transition. Moreover, the phase lags and coherence of J1727 show a significant excess at $\sim$0.4 Hz and $\sim$3 Hz – previously observed in other sources, with a different instrument (Méndez et al., 2024; König et al., 2024) – confirming a physical origin of the signal.
@@ 4.4 QPO energy dependence
In J1727, the QPO frequency does not depend on energy, it stays constant in the 0.3–2 keV compared to the 2–12 keV energy bands. At the QPO frequency ($\sim$2 Hz), the fractional rms continually increases with energy, due to the decreasing contribution of the disk at higher energies (as observed in other sources; Alabarta et al., 2022, 2025; Fogantini et al., 2025; Bellavita et al., 2025). Parallelly, the phase lags reach a minimum value around 1 keV for J1727, around 1.6 keV for MAXI J1348–630 (Alabarta et al., 2025) and around 1.5 keV for MAXI J1820+070 (Ma et al., 2023), staying constant $\gtrsim$4 keV. The U-shaped energy lags could be explained by a fraction of the cold seed photons from the accretion disk entering the extended corona in the horizontal direction, resulting in a significant drop in the phase lags (Liao et al., 2024; Alabarta et al., 2025; Bellavita et al., 2025). Furthermore, as the source hardens, Bellavita et al. (2025) show a more prominent drop in the phase lags.

### v2

@@ 3.2 The real and imaginary parts of the cross spectrum
Following Méndez et al. (2024), the real and imaginary parts of the CS were fit in a similar way as the PS. They were decomposed into a linear combination of Lorentzians multiplied by the phase-lag dependence using the relation
@@
with $C=\sqrt{AB}$, and the frequency-dependent phase lags between the two signals were defined as
@@
We explicitly treated the phase lags of the ith Lorentzian as a function of the parameters $p_{\textit{j,i}}$. From now on, we assume the phase lags of each Lorentzian to be constant with frequency, that is, $g_{i}(\nu;p_{j,i})=\Delta\phi_{\textit{i}}$, where $\Delta\phi_{\textit{i}}$ is a parameter of the model. This is the constant phase-lag model as defined by Méndez et al. (2024). The phase lags of J1727 are close to zero across the Fourier frequencies, and the amplitude of the imaginary part of the CS is therefore smaller than that of the real part. Consequently, the cross vectors can be rotated by 45° to have comparable real and imaginary parts. This generates more stable fits in Xspec and does not affect the best-fitting parameters (see Méndez et al., 2024). By multiplying the cross vector (decomposed into a real and imaginary part) by the rotation matrix, we therefore calculated the real and imaginary parts of the CS as $\mathrm{\Re[\cos{(\pi/4)}]\ -\ \Im[\sin{(\pi/4)}]}$ and $\mathrm{\Re[\sin{(\pi/4)}]\ +\ \Im[\cos{(\pi/4)}]}$, respectively. The central frequencies and FWHMs of the three Lorentzians in the CS are tied to those in the 0.3–2 keV band of the PS. The normalization of the real part of the CS is given by the parameter $C$, as previously defined. Table 1 shows the best-fitting parameters of the phase lags of the CS between 0.3–2 keV and 2–12 keV. The rotated real (green) and rotated imaginary (pink) parts of the CS we fit are shown in the right panel of Fig. 1. Below 1.5 Hz, the rotated real and imaginary parts of the CS overlap in the two energy bands for both observations. Above 1.5 Hz, the rotated imaginary part of the CS in the 2–12 keV energy range is larger than the rotated real part. A bump is observed in the rotated imaginary part of the CS at $\approx$3 Hz, and it is well fit by a narrow Lorentzian. In both observations, we identify a narrow Lorentzian in the PS and the CS with a larger imaginary component than in the real part. We call this QPO the imaginary QPO, as first expressed by Méndez et al. (2024). While the rms of the imaginary QPO stays more or less constant when the system hardens, the real part of the QPO increases.
Figure 2: Phase lags in radian (top) and the coherence function (bottom) for the second (left) and third (right) EPIC-PN observation of Swift J1727.8–1613. The continuous lines show the predicted model from the fits of the PS and CS represented in Fig. 1. The bottom panels show the residuals defined as (data-model)/error. The dotted vertical red line indicates the central frequency of the imaginary QPO for each observation.
@@ 3.3 The phase lags and coherence
The last column of Table 1 shows the best-fitting parameters for the phase lags. The phase lags of the three Lorentzians increase by $\approx$37%, $\approx$13%, and $\approx$19% (for L1, L2, and L3, respectively) between the second and third observation. From the PS and CS fits, we derived a model of the phase lags and the coherence, calculated in the same energy bands as the power and cross spectra. Fig. 2 displays the phase lags (top panel) and the coherence (bottom panel) for the two observations, together with the predicted model. Below 1 Hz, the phase lags are constant around $\approx$0 rad on average, with a slight excess at $\approx$0.2 Hz and $\approx$0.5 Hz that is significant at the 68% confidence level. Then, the phase lags reach a maximum value at the imaginary QPO frequency, marked by the vertical dotted red line, and stabilize to a higher constant value at $\approx$1.8 rad after this for both observations. In the second observation, however, the phase lags are narrower and are dominated by the uncertainties above 7 Hz for both observations.
While the phase lags increase at the central frequency of the imaginary QPO, the coherence drops. Below 1 Hz, the average coherence value is $\approx$1, it then drops to $\approx$0.8 at the frequency of the imaginary QPO and remains constant at $\approx$0.9 after the dip. Similarly to the phase lags, the residuals drop significantly at $\approx$0.5 Hz, aligning with a slight decrease in the coherence. These features suggest that an additional component contributes to the variability at low frequency. We did not add another Lorentzian to account for this component in our model, because we aimed to fit the prominent variable components with as few Lorentzians as possible.
@@ 3.4 The energy dependence of the QPO lags and rms amplitude
To analyze the QPO lags and rms amplitude energy dependence, we computed the fast Fourier transforms (FFTs) and CS for each narrow energy band relative to the full (0.3–12 keV) energy band at the imaginary QPO frequency. The data were then fit using the multi-Lorentzian model we defined above, and the central frequencies and FWHMs were fixed to the best-fit values listed in Table 1. Fig. 3 shows the energy dependence of the fractional rms and the phase lags.
@@
On the other hand, the phase lags reached a minimum value at 1 keV. In epoch XMM2, the phase lags decreased from $\approx$ –0.08 rad to $\approx$ –0.28 rad between 0.5 keV and 1 keV and increased again from $\approx$ –0.21 rad at 1.2 keV to $\approx$0.58 rad at 8 keV. In epoch XMM3, a similar behavior is observed with a decrease from $\approx$ –0.09 rad at 0.5 keV to a minimum value of $\approx$ –0.31 rad at 1 keV that increased above 1 keV.
@@ 4 Discussion and conclusions
In section 3, we showed that the XMM EPIC-PN data of the BH LMXB Swift J1727.8–1613 reveal a type-C QPO in the soft–to–hard transition. While the characteristic Lorentzians frequencies shift to lower values as the source hardened, the phase lags increased and the coherence function dropped at the QPO frequency. This supports again the possible cause–to–effect relation between the type-C QPO and the features observed in the lags and the coherence. An in-depth discussion is provided below.
@@ 4.1 The type-C QPO
We identified a narrow Lorentzian in the PS characterized by a higher rms amplitude in the imaginary part than in the real part of the CS, where the rms of the real counterpart increased between the second and the third observation. The origin of the imaginary QPO is still unclear: In Cyg X-1, König et al. (2024) suggested a beat between the two broad Lorentzians, that might create a third component causing the QPO, as is often observed in the PS, whereas Méndez et al. (2024) argued that it comes from an independent phenomenon. Nonetheless, the imaginary QPO of the source seems to correspond to the QPO observed in the PS.
@@
The QPO identified in J1727’s PS and CS has a central frequency at $\approx$2.6 Hz in the second observation that shifts to $\approx$2 Hz in the third observation. Furthermore, the total rms is $\approx$17% and $\approx$18% in the 0.3–2 keV energy band for the second and last observation, respectively, and it reaches $\approx$30% in the 2–12 keV energy band for both observations. Based on the brightness of J1727 and on high-rms LFQPO in the PS and CS, we classified the LFQPO that is observed in the soft–to–hard transition as a type-C QPO (Motta et al., 2011). In the rms–intensity diagram for black hole transients defined by Muñoz-Darias et al. (2011), the source would be close to the hard line, which is defined as a linear relation between the absolute rms and the count rate observed in the canonical low-hard state. Weak type-C QPOs are generally observed at the hard line in the hard state and migrate above the line as the rms increases and reaches 30%, and they progressively leave the hard line. The $Q$ factor values for J1727 are lower than those usually observed for type-C QPOs ($\approx$7–10), however. Low $Q$ factors in type-C QPOs were also observed in MAXI J1348–630 (Alabarta et al., 2022) and MAXI J1820+070 (Ma et al., 2023), where they ranged between 0.2 and 7.
@@
On the other hand, as the system hardened, the QPO moved to lower frequencies. This frequency shift (probably induced by the Comptonizing plasma that dominates the emission in the hard state) has previously been observed in other XBs (MAXI J1820+070; Méndez et al., 2024; Bellavita et al., 2025, and Cyg X-1; König et al., 2024; Fogantini et al., 2025). Furthermore, during the soft–to–hard transition of MAXI J1348–630, Alabarta et al. (2025) interpreted the shift in the QPO frequency to lower frequencies as indicative of coronal expansion. In the case of GRS 1915+105, Méndez et al. (2022) and García et al. (2022) explained the frequency shift by a perpendicular contraction of the corona with respect to the accretion disk. Thus, the frequency shift suggests a geometric change in the corona.
@@
Additionally, observed in the soft–to–hard transition, the type-C QPO can be attributed to the expansion of the truncated inner radius of the disk or to feedback processes from the corona (Bellavita et al., 2022). The Lense-Thirring precession, in which the inner accretion flow precesses with respect to the thermal disk (see Ingram et al., 2009), has also been proposed as the origin of type-C QPOs in XBs. This was supported by studies that reported a tight anticorrelation between the outer radius of the Comptonizing region and $\mathrm{\nu_{QPO}}$ (Kubota et al., 2024). The analysis of GRS 1915+105 by Nathan et al. (2022) revealed an inner disk radius that was too small and a reprocessing time that was too long to support this model, however.
@@
In addition to the type-C QPO, the broadband Lorentzians L1 and L3 similarly shift to lower frequencies as the source hardens. In the 0.3–2 keV energy band, however, their rms amplitude increases as the source transitions, unlike that of the type-C QPO. This behavior was previously observed in GX 339-4 (Muñoz-Darias et al., 2011). In the 2–12 keV hard band, the rms increases for L3 for a harder emission, but it decreases for L1. Thus, as suggested by Bellavita et al. (2025), the two broad Lorentzians seem to be independent from each other and might come from variability processes that differ from the QPO. König et al. (2024) supported this assumption by proposing that L1 comes from the accretion disk and is modulated by the Comptonization and that L3 is associated with the Comptonized medium. On the other hand, Wilkinson & Uttley (2009) linked the L1 variability to an unstable accretion flow in the disk (although they reached this conclusion by fitting the covariance spectrum with the same model as the time-averaged spectrum, which only holds when the overall normalization of the model alone varies).
4.2 Phase-lag jump
@@
At the QPO frequency, the phase lags reach a maximum, also called the cliff. This was previously observed in other XBs (Méndez et al., 2024; Bellavita et al., 2025). Before and after the QPO frequency, however, the phase lags remain constant at $\approx$0 rad and $\approx$0.18 rad, respectively, for both observations, with a sharper cliff in the third observation. In the frequency range in which one Lorentzian dominates the PS, the phase lags remained constant on average, where each Lorentzian had its own lags, potentially caused by separate resonances (Méndez et al., 2024). At the intersection of the two Lorentzians, in contrats, the phase lags transition from one constant value to the next. Nowak et al. (1999b) attributed this difference to the total phase lags at each Fourier frequency being equal to the average phase lags of the individual Lorentzians, weighted by the product of their amplitudes in the two energy bands at which they were calculated (defined as the small-angle approximation of equation 9 in Méndez et al., 2024). As a result, the phase lags are constant over a given frequency range when the Fourier amplitudes of one Lorentzian dominate.
@@
Moreover, no cliff is observed in the hard-energy bands. Similarly observed in the rising phase of the outburst of J1727, Yu et al. (2024) did not find a phase-lag jump in the 1–10 keV and 25–150 keV energy bands, but still detected it in the soft energy bands, 1–2 keV and 2–10 keV. This might suggest that soft energies are essential to observe the phase-lag jump, regardless of the transition direction. This similarly indicates that the QPO originated from variations in a soft component that is linked to the accretion disk. Furthermore, as the source hardens, the Comptonization-driven variability from the corona intensifies and might introduce additional variability in the accretion disk.
@@ 4.3 Coherence drop
The phase-lag jump and the coherence drop seem to be generated overall by low-energy variations from a soft component in the accretion disk at the type-C QPO frequency, induced by variations in the corona geometry (Méndez et al., 2022; Alabarta et al., 2025). Only observed in the decaying phase of the outburst (Méndez et al., 2024; Alabarta et al., 2025; König et al., 2024; Bellavita et al., 2025; Fogantini et al., 2025), these features are significant at low X-ray flux in the soft–to–hard transition. Moreover, the phase lags and coherence of J1727 show a significant excess at $\approx$0.4 Hz and $\approx$3 Hz that was previously observed in other sources with a different instrument (Méndez et al., 2024; König et al., 2024; Rout et al., 2025). This confirms the physical origin of the signal.
@@ 4.4 QPO energy dependence
[…]